OpenStudy (dmezzullo):

Solve: 4 − x = 19 15 −15 21 −21

OpenStudy (anonymous):

add -4 to both sides of equation -4+4-x=19-4 so it becomes -x=15 now Multiply with minus on both sides to get exact answer ...